ToxicaDennisIce:

Vanessa owed her friend $24. She paid back $8. How much more does Vanessa need to pay before her account is at zero? A –$16 B –$8 C $8 D $16

yesno:

24-8 is what? count..

Hoodmemes:

How she got 16 \[24-8= 16\] And yes D is correct.

yesno:

okay thanks.

yesno:

n yea

yesno:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @supie \(\color{#0cbb34}{\text{Originally Posted by}}\) @Hoodmemes How she got 16 \[24-8= 16\] And yes D is correct. \(\color{#0cbb34}{\text{End of Quote}}\) Yeah, like that. \(\color{#0cbb34}{\text{End of Quote}}\) wait but why is there "A $16 " then "D $16" ? confused.

Hoodmemes:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @yesno \(\color{#0cbb34}{\text{Originally Posted by}}\) @supie \(\color{#0cbb34}{\text{Originally Posted by}}\) @Hoodmemes How she got 16 \[24-8= 16\] And yes D is correct. \(\color{#0cbb34}{\text{End of Quote}}\) Yeah, like that. \(\color{#0cbb34}{\text{End of Quote}}\) wait but why is there "A $16 " then "D $16" ? confused. \(\color{#0cbb34}{\text{End of Quote}}\) thats -16

supie:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @yesno \(\color{#0cbb34}{\text{Originally Posted by}}\) @supie \(\color{#0cbb34}{\text{Originally Posted by}}\) @Hoodmemes How she got 16 \[24-8= 16\] And yes D is correct. \(\color{#0cbb34}{\text{End of Quote}}\) Yeah, like that. \(\color{#0cbb34}{\text{End of Quote}}\) wait but why is there "A $16 " then "D $16" ? confused. \(\color{#0cbb34}{\text{End of Quote}}\) Option A says -16(negative 16) and Option D says +16(regular 16, positive 16)

yesno:

"A –$16 B –$8 C $8 D $16" same answers- twice..

Hoodmemes:

there is a minus sign

yesno:

i dont se it lol.

supie:

Option A says -16(negative 16) and Option D says +16(regular 16, positive 16)